A right cylindrical can is to hold 1 L. of water. What should be the height of the can if the radius is 12 cm? Use 3.14159.
The Height should be approximatelym _______ cm. (nearest hundredth)

Answer by edjones(8007)

1 L = 1000 cm^3
pi*r^2*h=V(cylinder)
3.14159*12^2*h=1000
452.38896*h=1000
h=2.21 cm
.
Ed